Cindy earns a 15% commission on all sales. On Saturday, she sold $985.40 worth of merchandise.
Answer by tutorcecilia(2152)

Use the formula for percentages:
Part/Whole=%/100
x/985.40 = 15%/100 [Solve for x]
(x)(100) = (985.40)(15%)
100x = 14781
100x/100 = 985.40/100
x = $147.81 was her earned commission.